Reports 1-1 of 1 Clear search Modify search
AdV-FLT (Installation)
capocasa, vardaro - 20:10 Wednesday 16 June 2021 (52173) Print this report
Comment to Baffle installation on the END mirror (52171)

FCEM marionette has been properly oriented looking at the back reflection from the open flange. We did the adjustment with open loop.

  • TX has been adjusted moving the marionette TX motor (number 5)
  • TY has been adjusted adding a DC on the coils (0.15 V)

Then we centered oplev PSD and closed the loop. We needed to further adjust the loop set-points to find the flashes. (TX = -100, TY= -500)

We have also realigned PSD2 (TZ), PSD3 (MIR_TY, MIR_TX) and PSD 4 (MIR_Z) 

Config file has been updated with the new values.

Search Help